OpenStudy (mikezack123):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = −1/2x − 2 y = 2x − 2 y = −1/2x − 7

OpenStudy (anonymous):

Parallel=same slope

OpenStudy (anonymous):

So plug in the same slop into point slope with the point (5,3) like before and you'll get one of these answers^^^^

OpenStudy (anonymous):

do you get what he said? or do u want me to explain it?

OpenStudy (mikezack123):

i want the answer!

OpenStudy (texaschic101):

we are not allowed to just give the answers...we can get suspended...believe me, I know from past experience

OpenStudy (texaschic101):

chris said he could explain it...you should listen to him and learn

OpenStudy (anonymous):

alright. parralel lines have the same slope. ok

OpenStudy (anonymous):

so any parralel line has the form

OpenStudy (anonymous):

but this particular line must travel through a particular point. this point being (5, 3)

OpenStudy (anonymous):

i.e x=5 y=3

OpenStudy (mikezack123):

okay im keeping up...

OpenStudy (anonymous):

so what do u think we should do with x=5 y=3?

OpenStudy (texaschic101):

it is going to have the same slope....so what is the slope of the parallel line your looking for ?

OpenStudy (texaschic101):

y = mx + b....the slope being m Look at your equation...y = 2x + 2...what is the slope ?

OpenStudy (mikezack123):

Got it the anwer is A!

OpenStudy (texaschic101):

let me check...hold on

OpenStudy (anonymous):

correct

OpenStudy (mikezack123):

woo hoo!

OpenStudy (texaschic101):

y = mx + b using slope 2 and points (5,3) 3 = 2(5) + b 3 = 10 + b 3 - 10 = b -7 = b your new equation is : y = 2x - 7 you are correct :)

OpenStudy (mikezack123):

i dont know who to give my medal to... im sorry you all helped! :( sooooo, first to respond to this question with an @mikezack123 gets the medal

OpenStudy (mikezack123):

only IF you helped

OpenStudy (texaschic101):

he should get it.....I agree

OpenStudy (anonymous):

no fuss mate

OpenStudy (texaschic101):

no fuss...all is good :)

OpenStudy (mikezack123):

so, Thanks to the both of you! ill fan you both.

OpenStudy (texaschic101):

thats what we are here for :)